OpenStudy (anonymous):

solve for x 5/2x + 1/4x = 7/4 + x thank you

OpenStudy (anonymous):

I got two different answers when I went to check my work so maybe somebody could correct me?

OpenStudy (anonymous):

10/4x +1/4x=7/4+4/4x 11/4x=7/4 + 4/4x (11/4x-4/4x)=7/4x 7/4x=7/4 x=28/28 or x=1 is my first answer

OpenStudy (anonymous):

this is my second answer 20/4x + 1/4x = 7/4 + x 21/4x = 7/4 + x 4.25x = 7/4 x = (7/4)/(4.25) 21/4 - 1 21/4 - 4/4 = 17/4 7/4 * 4/17 = 28/68

OpenStudy (mertsj):

But in you second problem you have that 5/2 is 20/4 and that is not true

OpenStudy (mertsj):

You could cheeck you answer using substitution if you wanted to know which one is correct.

OpenStudy (anonymous):

okay so its probably my first answer! i have to get off though so you're welcome and good luck :)

OpenStudy (anonymous):

thank you for trying i really appreciate it :)

OpenStudy (anonymous):

no problem! :)